OpenStudy (anonymous):

Will give medal What is the sum of this infinite geometric series? 2 + 1 + 1/2+1/4+1/8…

zepdrix (zepdrix):

Ignore the 2 for a sec, that's kind of messing things up. 1 + 1/2+1/4+1/8… So we have a common ratio of 1/2 yes? We can write this as:\[\Large\rm \sum_{n=0}^{\infty}\left(\frac{1}{2}\right)^n\] And if you recall, the convergent geometric series can be written as:\[\Large\rm \sum_{n=0}^{\infty}\left(r\right)^n=\frac{1}{1-r}\]

zepdrix (zepdrix):

So then dealing with the 2, I guess we have: \[\Large\rm 2+\color{orangered}{\Large\rm \sum_{n=0}^{\infty}\left(\frac{1}{2}\right)^n}\]Which is:\[\Large\rm 2+\color{orangered}{\frac{1}{1-\frac{1}{2}}}\]Yes? :o

OpenStudy (anonymous):

So, would the answer be 4? Or am I missing something?

zepdrix (zepdrix):

Ooo I maybe messed up the numerator there >.< lemme check...

zepdrix (zepdrix):

nonono I think that's right.. Yah 4 sounds right. :)